# On Continuous Optimization for Constraint Satisfaction Problems

### v1

@@ Appendix B Benchmarks and Encodings
Task Scheduling Problem
@@
Graph Coloring Problem with Hashing Queries
@@
For any $h\in{}H$, as inspired by the Fourier expansion of XOR constraints (Kyrillidis et al. 2020), the circuit-output probability of the constraint $c\in{}C_{2}$ can be written as
@@
After the above relaxations, the main computation of FourierCSP can be amenable for parallelization using Jax (Bradbury et al. 2018).
@@
Appendix C Convergence Analysis
We selected 36 test cases to analyze convergence behavior. We choose $n\in\{4,8,16\}$ and $m\in\{2,3,4\}$ to generate the test cases, where $n$ is the number of variables and $m$ is the number of states of each variable. The constraint expression consists of the variables and the connectives $\wedge,\vee,\oplus,<,>$, along with parentheses for grouping. Representative examples for $n=4$ and $n=8$ are shown in Fig.3 and Fig.4, respectively.
For all cases, the projected gradient aims to maximize the $m_{td}$ of the true terminal node, or equivalently, minimize $1-m_{td}[true]$. The gradient updates are performed using jaxopt.ProjectedGradient with default parameters. Each run is initialized at the uniform point $\left(\frac{1}{m},\cdots,\frac{1}{m}\right)^{n}$. We record the number of iterations required for convergence and identify whether the final solution lies on the boundary or at a corner of the domain. The results are shown in Table 1,  2, and  3. In every test case, the projected gradient converges to a point $\mathbb{P}[true]=1$. Moreover, each convergent point lies at a corner of the domain, i.e., every variable can be deterministically rounded to a discrete state.
Figure 3: Decision diagrams of the test cases with 4 variables. Left: $c_{1},c_{2},c_{3}$; Mid: $c_{13},c_{14},c_{15}$; Right: $c_{25},c_{26},c_{27}$.
Figure 4: Decision diagrams of the test cases with 8 variables. Left: $c_{5},c_{6},c_{7}$; Mid: $c_{17},c_{18},c_{19}$; Right: $c_{29},c_{30},c_{31}$.
#Variable Name Expression #Iter. On Boundary On Corner
4 $c_{1}$ $x_{1}\land x_{2}\land x_{3}\land x_{4}$ 3 $x_{1},x_{2},x_{3},x_{4}=\{1,1,1,1\}$
$c_{2}$ $x_{1}\lor x_{2}\oplus x_{3}\land x_{4}$ 4 $x_{1},x_{2},x_{3},x_{4}=\{1,1,0,1\}$
$c_{3}$ $x_{1}\land x_{2}\land x_{3}\lor x_{4}$ 2 $x_{4}=\{1\}$
$c_{4}$ $c_{2}\wedge c_{3}$ 4 $x_{1},x_{2},x_{3},x_{4}=\{1,1,0,1\}$
8 $c_{5}$ $x_{1}\land\cdots\land x_{8}$ 6 $x_{1},\ldots,x_{8}=\{1,\ldots,1\}$
$c_{6}$
$x_{1}\oplus x_{2}\land x_{3}\lor x_{4}$
$\land x_{5}\land x_{6}\lor x_{7}\oplus x_{8}$
5 $x_{7},x_{8}=\{1,0\}$
$c_{7}$
$x_{1}\land x_{2}\oplus x_{3}\land x_{4}$
$\lor x_{5}\land x_{6}\oplus x_{7}\land x_{8}$
4 $x_{6},x_{7},x_{8}=\{0,1,1\}$
$c_{8}$ $c_{6}\wedge c_{7}$ 8 $x_{1},\ldots,x_{8}=\{1,1,0,0,1,1,0,1\}$
16 $c_{9}$
$x_{1}\land x_{2}\land\cdots\land x_{16}$
12
$x_{1},\ldots,x_{16}=\{1,\ldots,1\}$
$c_{10}$
$x_{1}\land x_{2}\land x_{3}\lor x_{4}\lor x_{5}\oplus x_{6}$
$\land x_{7}\oplus x_{8}\lor x_{9}\land x_{10}\oplus x_{11}$
$\land x_{12}\lor x_{13}\land x_{14}\lor x_{15}\lor x_{16}$
21 $x_{16}=\{1\}$
$c_{11}$
$x_{1}\oplus x_{2}\land x_{3}\lor x_{4}\land x_{5}\oplus x_{6}$
$\land x_{7}\oplus x_{8}\lor x_{9}\land x_{10}\oplus x_{11}$
$\lor x_{12}\land x_{13}\land x_{14}\oplus x_{15}\land x_{16}$
9 $x_{13},x_{14},x_{15},x_{16}=\{0,0,1,1\}$
$c_{12}$ ${c_{10}}\wedge{c_{11}}$ 7 $x_{13},x_{15},x_{16}=\{0,1,1\}$
Table 1: Convergence analysis of CSP formula where each variable has 2 states.
#Variable Name Expression #Iter On Boundary On Corner
4 $c_{13}$ $x_{1}\land x_{2}\land x_{3}\land x_{4}=1$ 4 $x_{1},x_{2},x_{3},x_{4}=\{1,1,1,1\}$
$c_{14}$ $(x_{1}<x_{2})\oplus(x_{3}>x_{4})=1$ 14 $x_{1},x_{2},x_{3},x_{4}=\{2,0,2,0\}$
$c_{15}$ $x_{1}\land(x_{2}<x_{3})\land x_{4}=1$ 4 $x_{1},x_{2},x_{3},x_{4}=\{1,0,2,1\}$
$c_{16}$ $c_{14}\wedge c_{15}$ 3 $x_{1},x_{2},x_{3},x_{4}=\{1,0,2,1\}$
8 $c_{17}$ $x_{1}\land x_{2}\land\cdots\land x_{8}=1$ 9 $x_{1},\ldots,x_{8}=\{1,\ldots,1\}$
$c_{18}$
$\big(x_{1}\land x_{2}\land x_{3}\lor x_{4}\oplus$
$(x_{5}<x_{6})\land x_{7}\land x_{8}\big)=1$
12 $x_{5},x_{6}$ $x_{4},x_{7},x_{8}=\{1,1,1\}$
$c_{19}$
$\big(x_{1}\lor x_{2}\oplus(x_{3}<x_{4})\lor$
$x_{5}\land(x_{6}<x_{7})\oplus x_{8}\big)=1$
6 $x_{6},x_{7}$ $x_{8}=\{1\}$
$c_{20}$ $c_{18}\wedge c_{19}$ 7 $x_{1},x_{2},x_{3},x_{6}$ $x_{4},x_{5},x_{7},x_{8}=\{1,2,1,1\}$
16 $c_{21}$
$x_{1}\land x_{2}\land\cdots\land x_{16}=1$
108 $x_{1},\ldots,x_{16}=\{1,\ldots,1\}$
$c_{22}$
$\big((x_{1}<x_{2})\land(x_{3}>x_{4})$
$\oplus x_{5}\land(x_{6}<x_{7})\oplus x_{8}$
$\land x_{9}\land x_{10}\lor x_{11}\land x_{12}\oplus$
$x_{13}\lor x_{14}\oplus(x_{15}<x_{16})\big)=1$
139 $x_{15},x_{16}$ $x_{12},x_{13}=\{0,1\}$
$c_{23}$
$\big((x_{1}>x_{2})\land x_{3}\land(x_{4}<x_{5})$
$\lor x_{6}\oplus(x_{7}>x_{8})\land x_{9}$
$\lor x_{10}\lor(x_{11}>x_{12})\lor$
$(x_{13}<x_{14})\land(x_{15}<x_{16})\big)=1$
7 $x_{3},x_{6},x_{9}$ $x_{10},x_{15},x_{16}=\{1,0,2\}$
$c_{24}$ $c_{22}\wedge c_{23}$ 96 $x_{3},x_{6},x_{9}$ $x_{10},x_{12},x_{13},x_{14},x_{15},x_{16}=\{1,0,0,0,0,2\}$
Table 2: Convergence analysis of CSP formula where each variable has 3 states.
#Variable Name Expression #Iter. On Boundary On Corner
4 $c_{25}$ $x_{1}\land x_{2}\land x_{3}\land x_{4}=1$ 5 $x_{1},x_{2},x_{3},x_{4}=\{1,1,1,1\}$
$c_{26}$ $x_{1}\land(x_{2}<x_{3})\land x_{4}=1$ 5 $x_{1},x_{2},x_{3},x_{4}=\{1,1,2,1\}$
$c_{27}$ $x_{1}\land x_{2}\land(x_{3}>x_{4})=1$ 4 $x_{3},x_{4}$ $x_{1},x_{2}=\{1,1\}$
$c_{28}$ $c_{26}\wedge c_{27}$ 3 $x_{3}$ $x_{1},x_{2},x_{4}=\{1,1,1\}$
8 $c_{29}$ $x_{1}\land\cdots\land x_{8}=1$ 11 $x_{1},\ldots,x_{8}=\{1,\ldots,1\}$
$c_{30}$
$\big((x_{1}<x_{2})\land(x_{3}>x_{4})\land$
$x_{5}\lor(x_{6}\oplus x_{7})\land x_{8}\big)=1$
4 $x_{1},x_{2},x_{3},x_{4}$ $x_{5},x_{8}=\{1,1\}$
$c_{31}$
$\big(x_{1}\land x_{2}\oplus x_{3}\lor(x_{4}<x_{5})$
$\oplus x_{6}\land x_{7}\oplus x_{8}\big)=1$
3 $x_{7},x_{8}=\{0,1\}$
$c_{32}$ $c_{30}\wedge c_{31}$ 63 $x_{5}$ $x_{6},x_{7},x_{8}=\{1,0,1\}$
16 $c_{33}$
$x_{1}\land x_{2}\land\cdots\land x_{16}$
145
$x_{1},\ldots,x_{16}=\{1,\ldots,1\}$
$c_{34}$
$x_{1}\oplus(x_{2}>x_{3})\land x_{4}\lor x_{5}$
$\oplus(x_{6}>x_{7})\lor(x_{8}<x_{9})$
$\land x_{10}\land(x_{11}<x_{12})\oplus x_{13}$
$\lor(x_{14}>x_{15})\oplus x_{16}$
52 $x_{11},x_{12}$ $x_{13},x_{16}=\{1,0\}$
$c_{35}$
$\big((x_{1}<x_{2})\land(x_{3}<x_{4})$
$\land x_{5}\oplus x_{6}\land x_{7}\oplus(x_{8}>x_{9})$
$\land x_{10}\land x_{11}\lor x_{12}\lor$
$(x_{13}>x_{14})\land(x_{15}>x_{16})\big)=1$
31 $x_{11},x_{15},x_{16}$ $x_{10},x_{12}=\{0,1\}$
$c_{36}$ $c_{35}\wedge c_{36}$ 41 $x_{15}$ $x_{10},x_{11},x_{12},x_{13},x_{16}=\{0,0,1,1,0\}$
Table 3: Convergence analysis of CSP formula where each variable has 4 states.
Appendix D Parallel Decision Diagram Traversal
The decision diagram can be traversed in parallel. For example, MDD $c_{13}$ and $c_{15}$ can be represented as Listing 1 and Listing 2.
@@ Appendix D Parallel Decision Diagram Traversal
Listing 2: c15.mdd
For every MDD, we call the kernel functions to traverse one edge at a time. The corresponding top-down and bottom-up message propagation kernels are shown in Listing 3 and Listing 4. To ensure uniform execution across threads, shorter MDD traces are padded with zeros. The kernel functions are designed to detect such padded entries (e.g., when nid equals zero) and skip message propagation accordingly.

### v2

@@ Appendix B Benchmarks and Encodings
B.1 Task Scheduling Problem
@@
B.2 Graph Coloring Problem with Hashing Queries
@@
For any $h\in{}H$, as inspired by the Fourier expansion of XOR constraints (Kyrillidis et al., 2020), the circuit-output probability of the constraint $c\in{}C_{2}$ can be written as
@@
After the above relaxations, the main computation of FourierCSP can be amenable for parallelization using Jax (Bradbury et al., 2018).
@@
Appendix C Efficient Evaluation and Differentiation by Decision Diagrams
Theorem 3.6 reveals that CSP solving can be reduced to a continuous optimization problem. The performance of gradient-based optimization crucially depends on the efficiency of evaluating and differentiating the objective function. However, the complexity of evaluation and differentiation is in #P-hard, i.e., Eq. (1) has at most exponentially many terms, in general. Motivated by (Thornton and Nair, 1994), we reformulate Eq. (1) to Circuit-Output Probability (COP), which admits more tractable computation when a structured representation can encode the underlying constraint, i.e., decision diagrams.
Definition C.1 (Circuit-Output Probability).
Let $f_{c}$ be a discrete function over a set $X$ of finite domain variables. Let $P$ be the variable input probabilities. The circuit-output probability problem of $f_{c}$ w.r.t $P$, denoted by ${COP}_{c}(P)$, is the probability of $f_{c}$ outputs $1$ given the value of each variable independently sampled from a multimodal distribution, i.e.,
${COP}_{c}(P)=\sum_{X\in D}f_{c}(X)\prod_{i=1}^{n}p_{i,x_{i}}.$
Corollary C.2.
$\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[f_{c}(X)]={COP}_{c}(P).$
Proof of Corollary C.2.
$=\mathop{\mathbb{P}}_{X\sim\mathcal{S}_{P}}[f_{c}(X)=1]$
$={COP}_{c}(P).$
Prior work demonstrates that BDD can encode many Boolean constraints in compact size (Kyrillidis et al., 2021b; Sasao et al., 1996; Aavani, 2011). The multiple-valued extension of BDD is called a multi-valued decision diagram (MDD) (Bryant, 1995). The MDD can be versatile in representing many constraints through the translation from table constraints (Cheng and Yap, 2010) [^1] A table constraint is explicitly represented as the set of solutions or non-solutions for that constraint.. Similar to BDD, two MDDs can be combined through an Apply operation (Bryant, 1986; Perez and Régin, 2015), offering flexibility in encoding versatile constraints. The MDDs can be reduced by merging nodes that have the same set of outgoing neighbors associated with the same label (Cheng and Yap, 2010; Perez and Régin, 2015). Moreover, the size can be further reduced by sifting (Miller and Drechsler, 2003). A compact-size decision diagram means that the COP on it can be evaluated and differentiated efficiently.
C.1 Evaluation
Lemma C.3 (Top-Down Evaluation).
$m_{td}[v]=\mathbb{P}_{M,a}[v],$
$m_{td}[\textsc{true}]={COP}_{c}(P_{a}),$
i.e., Algorithm 2 returns the COP of $G$ in time $O(|V|)$.
Proof of Lemma C.3.
We prove by induction on the MDD $G(V,E)$. Let $A\in\tilde{D}$ and $B\in{}D$.
  • Inductive step: For each non-root node $v$ of $V$, let $par_{i}(v)$ be ths set of parent nodes of $v$ with an edge labeled by $i$. After Algorithm 2 terminate, we have
    $m_{td}[v]=\sum_{u\in par_{i}(v)}a_{id(u),i}\cdot m_{td}[u],$
    $m_{td}[v]$ $=\sum_{u\in par_{i}(v)}\mathbb{P}[b_{id(u)}=i]\cdot\mathbb{P}[u|G,A]$
    $=\mathbb{P}[v|G,A]$
Algorithm 2 Top-Down Traversal of MDD
Output Circuit-output probability ${COP}_{c}(P)$.
1: Let $m_{td}[v]$ be the forward message of each node $v\in{}V$, where the root node is initialized as 1, otherwise 0.
2:for $v\in{}\texttt{sorted}(V)$ do
3:  for $u=\mathrm{child}(v)$ do
4:   $m_{td}[u]+=p_{\mathrm{id}(v),\mathrm{label}(v,u)}\cdot m_{td}[v]$
5:  end for
6:end for
7:return $m_{td}[true]$
Lemma C.4 (Bottom-Up Evaluation).
$m_{bu}[v]={COP}_{c,v}(P),$
i.e., Algorithm 3 returns the COP of $G$ and in time $O(|V|)$.
Proof of Lemma C.4.
We prove this by structural induction on MDD $G(V,E)$. Let $A\in\tilde{D}$, we have
    $m_{bu}[v]=\sum_{i\in\mathrm{dom}(x_{id(v)})}a_{id(v),i}\cdot m_{bu}[v.i].$
    By the inductive hypothesis, we have
    $m_{bu}[v]$ $=\sum_{i=0}^{k}a_{id(v),i}\cdot{COP}_{c,v.i}(A)$
    $=\sum_{i=0}^{k}\mathbb{P}[b_{id(v)}=i]\cdot{COP}_{c,v.i}(A)$
    $={COP}_{c,v}(A)$
    where ${COP}_{c,v}$ denotes the circuit-output probability of subgraph of $G$ rooted at node $v$.
Algorithm 3 Bottom-Up Traversal of MDD
Output Circuit-output probability ${COP}_{c}(P)$.
2:for $v\in{}\texttt{sorted}(V,\texttt{reverse=True})$ do
3:  for $u\in\mathrm{child}(v)$ do
4:   $i=\mathrm{id}(v)$
5:   $m_{bu}[v]+=p_{\mathrm{id}(v),\mathrm{label}(v,u)}\cdot m_{bu}[u]$
6:  end for
7:end for
8:return $m_{bu}[root]$
C.2 Differentiation
The main idea of differentiation is from probabilistic inference (Pearl, 1988; Shafer and Shenoy, 1990), which traverses the graph model top-down and bottom-up. The following theorem shows that the partial derivative of COP can be obtained by top-down and bottom-up messages from Algorithm 2 and 3.
Theorem C.5 (Derivatives).
$\frac{\partial{COP}_{c}(P)}{\partial p_{i,j}}=\sum_{v\in\{u:id(u)=i\}}m_{td}[v]\cdot m_{bu}[v.j]$ (13)
Proof of Theorem C.5.
Given an MDD $G(V,E)$, expand all terminal nodes such that each can only have one parent node. The expanded MDD is called $\mathcal{G}$, which has a true terminal set $T$ and a false terminal set $F$. We have the following by repeating the proof of Lemma C.3 and C.4 on $\mathcal{G}$.
${COP}_{c}(A)=\sum_{v\in{}T}\mathbb{P}[v|A,\mathcal{G}]=\sum_{v\in{}T}m_{td}^{\prime}(v)$
For any non-terminal node $v\notin(T\cup F)$,
$m_{td}[v]$ $=m_{td}^{\prime}[v],$
$m_{bu}[v]$ $=m_{bu}^{\prime}[v].$
For any true terminal node $v\in{T}$ and false terminal node $u\in{F}$,
$m_{bu}^{\prime}[v]$ $=1,$
$m_{bu}^{\prime}[u]$ $=0.$
Now, we are ready to prove Theorem C.5.
${COP}_{c}(A)$ $=\mathbb{P}[f_{c}=1|A,\mathcal{G}]$
$=\sum_{v\in\{u:id(u)=i\}}\mathbb{P}[v|A,\mathcal{G}]\cdot\mathbb{P}[f_{c}=1|v,A,\mathcal{G}]$
$=\sum_{v\in\{u:id(u)=i\}}m_{td}^{\prime}[v]\cdot m_{bu}^{\prime}[v]$
$=\sum_{v\in\{u:id(u)=i\}}m_{td}[v]\cdot m_{bu}[v]$
$=\sum_{v\in\{u:id(u)=i\}}m_{td}[v]\cdot\sum_{j=0}^{k}a_{i,j}m_{bu}[v.j]$
Hence we have
$\frac{\partial{COP}_{c}(P)}{\partial p_{i,j}}=\sum_{v\in\{u:id(u)=i\}}m_{td}[v]\cdot m_{bu}[v.j]$
C.3 Parallel Decision Diagram Traversal
The decision diagram can be traversed in parallel. For example, MDDs can be represented as Listing LABEL:lst:c13 and Listing LABEL:lst:c15.
@@ C.3 Parallel Decision Diagram Traversal
Listing 2: $x_{1}\wedge(x_{2}<x_{3})\wedge x_{4}$
For every MDD, we call the kernel functions to traverse one edge at a time. The corresponding top-down and bottom-up message propagation kernels are shown in Listing LABEL:lst:cuda_forward and Listing LABEL:lst:cuda_backward. To ensure uniform execution across threads, shorter MDD traces are padded with zeros. The kernel functions are designed to detect such padded entries (e.g., when nid equals zero) and skip message propagation accordingly.